An aircraft flew a distance of 3800 Km from Singapore to Perth in Australia at an average speed of v Km/h.
(a) Write down an expression in terms of v, for the time taken in hours for the journey.
The aircraft returned by the same route at anaverage speed of (v+50) Km/h.
(b) Write down an expression in terms of x, for the time taken in hours for the return journey.             
(c) Given that the difference in time between the two journeys is 20 minutes, form an equation in v and show that it reduces to V²+50v-570000=0

Answer by mananth(16946)

An aircraft flew a distance of 3800 Km from Singapore to Perth in Australia at an average speed of v Km/h.

(a) Write down an expression in terms of v, for the time taken in hours for the journey.
t=3800/v
The aircraft returned by the same route at anaverage speed of (v+50) Km/h.


(b) Write down an expression in terms of x, for the time taken in hours for the return journey.             
t=3800/(v+50)
(c) Given that the difference in time between the two journeys is 20 minutes, form an equation in v and show that it reduces to V²+50v-570000=0
3800/v -3800/(v+50)=1/3
(3800(v+50)-3800v)/v(v+50)= 1/3
190000v+19000-190000v = (v^2+50v)/3
570000 =v^2+50v
v^2+50v-570000=0